Write a code in matlab for AWGN () function. (please comment each line of the code to know what each line does)
Answer #1


f u n c ti o n [y,n] = add_awgn_noise(x,SNR_dB)
%[y,n]=awgn_noise(x,SNR) adds AWGN noise vector to signal ’x’ to generate a
%resulting signal vector y of specified SNR in dB. It also returns the
%noise vector ’n’ that is added to the signal x
L=l e n g t h(x);
SNR = 10ˆ(SNR_dB/10); %SNR to linear scale
Esym=sum(abs(x).ˆ2)/(L); %Calculate actual symbol energy
N0=Esym/SNR; %Find the noise spectral density
i f ( i s r e a l (x)),
noiseSigma = s q r t(N0);%Standard deviation for AWGN Noise when x is real
n = noiseSigma*randn(1,L);%computed noise
e l s e
noiseSigma=s q r t(N0/2);%Standard deviation for AWGN Noise when x is complex
n = noiseSigma*(randn(1,L)+1i*randn(1,L));%computed noise
end
y = x + n; %received signal
end